According to the ideal gas law, a 0.9249 mol sample of xenon gas in a 1.135 L container at 269.1 K should exert a pressure of 17.99 atm. By what percent does the pressure calculated using the van der Waals equation differ from the ideal pressure? For xe gas, a 34.194 L2atm/mol2 and b 5.105x102 L/mol Hint: difference 100x(P ideal van der Waals P ideal
Answer #1

Van der Waals equation-

\left ( P+\frac{an^{2}}{V^{2}} \right )(V-nb) = nRT

substituting the values, we get:

\left ( P+\frac{(4.194L^{2}atm/mol^{2})(0.9249mol)^{2}}{(1.135L)^{2}} \right )\left (1.135L-(0.9249mol)(5.105\times 10^{-2}L/mol) \right )

= (0.9249 mol)(0.0821L.atm/K.mol)(269.1K)

or, (P + 2.785 atm)(1.088 L) = 20.43 L.atm

(P + 2.785 atm) = 18.78 atm

P = 15.99 atm

Therefore, Van der Waals pressure = 15.99 atm

Ideal gas pressure = 17.99 atm

Pideal - PVanderWaals = 17.99 - 15.99 = 2.00 atm

% difference = (2.00/17.99)x100 =11.11 %